OpenStudy (anonymous):

Ivan, who is 1.72 m tall, wishes to find the height of a tree. He walks 22.47 m from the base of the tree along the shadow of the tree until his head is in a position where the tip of his shadow exactly overlaps the end of the tree top's shadow. He is now 5.03 m from the end of the shadows. How tall is the tree? Round to the nearest hundredth.

OpenStudy (nurali):

Total length of the shadow of tree is, 22.47m [as he walks from the base of tree in the shadow] + 5.03m [as at that point the shadow of tree and boy overlaps each other and is same]. = 27.5m A boy of height 1.72 casts a shadow 5.03m tall so, 1.72/5.03 = x/27.5 x=9.4